OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Write the equation for the vertical line that contains point E(10, -3).

OpenStudy (anonymous):

All vertica lines have the equation x = b, where b is any real number. Your x-value is 10, so the equation would be x = 10.

OpenStudy (anonymous):

oMGOSH thanks so much xx

OpenStudy (anonymous):

can you help with just one more?

OpenStudy (anonymous):



OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

Write the equation for the vertical line that contains point E(10, -3). Write it in Slope Intercept Form

OpenStudy (anonymous):

The vertical line, as we just said, is x = 10. It cannot be written in slope-intercept form. Slope-intercept form is y = mx + b. We cannot write x = 10 in that form.

OpenStudy (anonymous):

im sorry wrong question one sec i worte the wrong ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is an equation for the line that passes through points (-1, -4) and (1, 4)

OpenStudy (anonymous):

in slope intercept form

OpenStudy (anonymous):

First you need to find the slope of the line passing through those two points. Slope = difference in y-values/difference in x-values

OpenStudy (anonymous):

slope = (4 - (-4))/(1 - (-1)) = 8/2 = 4 So we know the slope, m, is 4.

OpenStudy (anonymous):

We want to write it in slope-intercept form, that is in the form y = mx + b. m = slope b = y-intercept., All what we need to do, is to find, b, the y-intercept. That's pretty easy. We have y = 4x + b (becuase the slope is 4).

OpenStudy (anonymous):

Now, take any of the two points you gave, say (1,-1), and plug it into the equation y = 4x + b, and we will solve for b. So, -1 = 4(1) + b - 1 = 4 + b b = - 5 So the equation will be y = 4x - 5.

OpenStudy (anonymous):

Thanks so much you don't know how much you have helped. You are better than my math teacher lol . Thank you!

OpenStudy (anonymous):

Welcome.